OpenStudy (anonymous):

Find the value of the expression: -6.4 + (-3.1)

OpenStudy (yrelhan4):

-->the expression = -6.4 - 3.1 = -8.5

OpenStudy (anonymous):

I got -3.3 but it was wrong

OpenStudy (yrelhan4):

a + (-b) = a-b

OpenStudy (mathstudent55):

To add two numbers with the same sign, poth positive or both negative, drop the signs and add the numbers, then give them the sign they had. -6.4 + (-3.1) You are adding two negative numbers, so drop the negatives and add them 6.4 + 3.1 = 9.5 Since both original numbers are negative, the result is negative: -6.4 + (-3.1) = -9.5

OpenStudy (anonymous):

so -9.5?

OpenStudy (yrelhan4):

oh crap! yes..

OpenStudy (anonymous):

Uhh?? lol